Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by BTGmoderatorRO » Sat Nov 11, 2017 8:30 am
The probability that event M will not occur is 0.8 and the probability that event R will not occur is 0.6. If events M and R cannot both occur, which of the following is the probability that either event M or event R will occur?

A) 1/5
B) 2/5
C) 3/5
D) 4/5
E) 12/25

OA is c

I solved this question twice and I got two answers, can someone confirm if it is C or A
Join the discussion
Source: — Problem Solving |

by EconomistGMATTutor » Sun Nov 12, 2017 9:17 am
Hello Roland2rule.

The question tell us that the probability that event M will NOT occur is 0.8, it implies that the probability that event M will OCCUR is 0.2.

Also, as the probability that event R will NOT occur is 0.6, then the probability that event R will OCCUR is 0.4.

Finally, it is asked what is the probability that either event M or event R will occur, so, the probability is the sum of the two probabilities calculated before. That is to say,
$$P\left(M\right)+P\left(R\right)=0.2+0.4\ =\ 0.6\ =\ \frac{3}{5}.$$

So, the correct answer is C .

I hope this can help you.

Feel free to ask me again if you have a doubt.
GMAT Prep From The Economist
We offer 70+ point score improvement money back guarantee.
Our average student improves 98 points.

Image
Join the discussion

by Scott@TargetTestPrep » Sat Oct 26, 2019 7:38 am
BTGmoderatorRO wrote:The probability that event M will not occur is 0.8 and the probability that event R will not occur is 0.6. If events M and R cannot both occur, which of the following is the probability that either event M or event R will occur?

A) 1/5
B) 2/5
C) 3/5
D) 4/5
E) 12/25

OA is c

I solved this question twice and I got two answers, can someone confirm if it is C or A

We are given that the probability that event M will not occur is 0.8 and the probability that event R will not occur is 0.6, and that events M and R cannot both occur.

We need to determine the probability that either event M or event R will occur.

The probability that event M will occur is 1 - 0.8 = 0.2 = 1/5

The probability that event R will occur is 1 - 0.6 = 0.4 = 2/5

Recall that the formula for the probability of event A or event B occurring is P(A or B) = P(A) + P(B) - P(A and B); therefore:

P(M or R) = P(M) + P(R) - P(M and R).

Since we are told that events M and R cannot both occur, that means P(M and R) = 0, and thus the probability that either event M or event R will occur is:

P(M or R) = 1/5 + 2/5 - 0 = 3/5.

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion